About Formatting a DMDW2000 Paper

Marie Dubois AFIP Ltd.

Hill Street 45, Manchester 67UR marie@afip.co.uk

John Q. Smith Computer Science Dept. Stanford, CA 94305 jqs@cs.stanford.edu

Peter Schmidt

Universität Mainz, Am Berg 10, 56652 Mainz peter@soci.org.de

Abstract

The Abstract paragraph should be indented 1/4 inch (2 pica) on both the left and right-hand margins. Abstract must be centered, bold, and in point size 12. Two line spaces precede the Abstract. The Abstract must be limited to one paragraph.

1 General Formatting Instructions

The format is in 2 columns with the overall line width of 6.75 inches (41 picas). The left margin is .9 inches (5.5 picas). Use 10 point type, with a vertical spacing of 11 points. Times Roman is the preferred typeface throughout.

Paper title is 16 point, Caps/lowercase, bold, centered. Subsequent pages should start at 1 inch (6 picas) from the top of the page.

Authors names are centered, initial caps; co-authors names, if used, are flush left and flush right.

Paragraphs are indented by 1 pica with no space between paragraphs.

2 First Level Heading

First level headings are all flush left, initial caps, bold and in point size 12. One line space before the first level heading and 1/2 line space after the first level heading.

2.1 Second Level Heading

Second level headings must be flush left, initial caps, bold and in point size 10. One line space before the second level heading and 1/2 line space after the second level heading.

2.1.1 Third Level Heading

Third level headings must be flush left, initial caps and bold. One line space before the third level heading and 1/2 line space after the third level heading.

Fourth Level Heading

Fourth level headings must be flush left, initial caps and roman type. One line space before the fourth level heading and 1/2 line space after the fourth level heading.

2.2 Citations In Text

Citations within the text should indicate the author’s last name and year[Knu73]. Reference style[Com79] should follow the style that you are used to using, as long as the citation style is consistent.

2.2.1 Footnotes

Indicate footnotes with a number^{1} in the text. Place the footnotes at the bottom of the page they appear on. Precede the footnote with a vertical rule of 2 inches (12 picas).

2.2.2 Figures

All artwork must be centered, neat, clean and legible. Do not use pencil or hand-drawn artwork. Figure number and caption always appear after the the figure. Place one line space before the figure, one line space before the figure caption and one line space after the figure caption. The figure caption is initial caps and each figure is numbered consecutively.

The copyright of this paper belongs to the paper’s authors. Permission to copy without fee all or part of this material is granted provided that the copies are not made or distributed for direct commercial advantage.

Proceedings of the International Workshop on Design and Management of Data Warehouses (DMDW’2000)

Stockholm, Sweden, June 5-6, 2000 (M. Jeusfeld, H. Shu, M. Staudt, G. Vossen, eds.)

http://sunsite.informatik.rwth-aachen.de/Publications/CEUR-WS/Vol-28/

1This is a sample footnote

Make sure that the figure caption does not get separated from the figure. Leave extra white space at the bottom of the page to avoid splitting the figure and figure caption.

Figure 1: Sample Figure Caption

2.2.3 Tables

All tables must be centered, neat, clean and legible. Do not use pencil or hand-drawn tables. Table number and title always appear before the table.

One line space before the table title, one line space after the table title and one line space after the table. The table title must be initial caps and each table numbered consecutively.

Table 1: Sample Table

A B 1
C D 2
E F 3

2.2.4 Acknowledgements

Use a third level heading for the acknowledgements. All acknowledgements go at the end of the paper.

2.2.5 Identification and Page Numbering

Insert the names of all authors at the bottom left of each page in 10 pt italics. The page number should appear at the bottom right preceded by the paper number. Please use the commands \def\shortauthors{J.Q. Smith, N. Coauthor} and \def\pagename{16-}.

2.2.6 References

Use a third level heading for the references. References follow the acknowledgements.

References

[Com79] D. Comer. The ubiquitous b-tree. Computing Surveys, 11(2):121–137, June 1979.

[Knu73] D. E. Knuth. The Art of Computer Programming – Volume 3 / Sorting and Searching. Addison-Wesley, 1973.